two cars start from the same place and at the same time but travel in opposite directions.after 6 hours, they are 540 metres apart. if one car travel 10 metres per hour slower than the other, what are their speed per hour?
Answer by mananth(16946)

speed of car A=x
speed of II car = x+10
They are moving away. so add up their speed
x+x+10=2x+10 m/h
distance = 540 m
time = 6 hours
speed = distance / time
2x+10 = 540/6
6(2x+10)=540
12x+60=540
12x=480
x=40 m/h
speed of car A = 40 m/h
speed of car II = 50 m/h